# 2Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 3Universität Heidelberg, Interdiszipliäres Zentrum für Wissenschaftliches Rechnen, D-69120 Heidelberg, Germany 4Dipartimento di Fisica e Astronomia Galileo Galilei, Università di Padova, Vicolo dell’Osservatorio 3, I–35122 Padova, Italy

### v1

@@
   Giuliano Iorio[Uncaptioned image]    Erika Korb
@@
Intermediate-mass black holes (IMBHs) are the link between stellar-mass and supermassive black holes. Gravitational waves have started unveiling a population of IMBHs in the $\sim 100-300\,{\rm M}_{\odot}$ range. Here, we investigate the formation of IMBHs from very massive stars (VMSs, $100\,{}{\rm M}_{\odot}$). We calculate new VMS models that account for the transition from optically thin to optically thick winds, and study how this enhanced mass loss affects IMBH formation and the black hole mass function at intermediate and high metallicity ($Z=10^{-4}-0.02$). We show that optically thick winds suppress the formation of IMBHs from direct VMS collapse at metallicities $Z>0.001$, one order of magnitude lower than predicted by previous models. Our models indicate that the stellar progenitors of GW231123 must have had a metallicity $Z<0.002$, if the primary black hole formed via direct VMS collapse.
@@ 1 Introduction
Intermediate-mass black holes (IMBHs), with masses $M_{\mathrm{BH}}\sim{10^{2}-10^{5}}$ M are the link between stellar-mass and supermassive black holes at the center of galaxies (Greene et al. 2020). In the last few years, gravitational waves have provided the first compelling evidence of their existence. The merger remnant of GW190521 represents the first IMBH ever detected via gravitational waves, with a mass of $142^{+28}_{-16}\,\mathrm{M_{\odot}}$ (Abbott et al. 2020, 2020). More recently, parameter estimation for the event GW231123 (The LIGO Scientific Collaboration et al. 2025c) reveals a black hole (BH) remnant with $m\sim 240\,{\rm M}_{\odot}$. Both BBH components of GW231123 have masses within or above the pair-instability mass gap (Heger & Woosley 2002; Woosley et al. 2007), with peculiarly high spins ($\chi\gtrsim 0.7$) that hint at the dynamical origin of this merger (Gerosa & Berti 2017; Fishbach et al. 2017; Antonini et al. 2019; Gerosa & Fishbach 2021; Mapelli et al. 2021; Antonini et al. 2025). The current fourth gravitational-wave transient catalog (The LIGO Scientific Collaboration et al. 2025a, b) already contains a number of other IMBH candidates, which is expected to increase with the next observing runs.
@@
At the same time, searches for IMBHs in the Milky Way globular clusters have mostly led to inconclusive results (Mezcua 2017), with only a few debated candidate BHs (Miller-Jones et al. 2012; Tremou et al. 2018; Baumgardt et al. 2019). One notable exception is represented by $\omega$ Centauri, which hosts an IMBH with mass $\geq{}8\times 10^{3}\,{\rm M}_{\odot}$, as revealed by proper motions based on over 20 years of HST data (Häberle et al. 2024).
@@
Possible IMBH formation pathways include the direct collapse of very massive stars (VMS), with masses above 100 M (Crowther et al. 2010; Bestenlehner et al. 2014; Vink et al. 2015; Costa et al. 2025; Shepherd et al. 2025), stellar collisions in dense young star clusters (Portegies Zwart & McMillan 2002; Freitag et al. 2006b, a; Mapelli 2016; Di Carlo et al. 2021; Torniamenti et al. 2022), and hierarchical BH mergers in dense and massive star clusters (Antonini et al. 2019, 2023; Mapelli et al. 2021; Rizzuto et al. 2021; Arca Sedda et al. 2023; Torniamenti et al. 2024).
@@
VMSs have been observed in star forming regions and young massive clusters (Vink et al. 2015), like the Arches cluster near the Galactic Center (Martins et al. 2008), NGC 3603 (Crowther et al. 2010), and R136 in the Large Magellanic Cloud (LMC, Crowther et al. 2010; Bestenlehner et al. 2014, 2020). The estimated initial masses of these objects are up to $\sim 300\,{\rm M}_{\odot}$ (Crowther et al. 2010). Also, most of them are observed as single stars (Crowther et al. 2010), which may hint that their formation is triggered or enhanced by stellar collisions during the earliest stellar phases (Portegies Zwart et al. 1999; Mapelli 2016). The possibility that VMSs form IMBHs, however, strongly depends on the amount of mass they lose during their life (Vink 2018).
@@
At solar metallicity, VMSs can lose a large fraction of their initial mass already during the main sequence, and leave relatively low-mass compact objects (<30<30 M, Belczynski et al. 2010; Mapelli et al. 2013; Romagnolo et al. 2024). At low metallicity, these stars could preserve enough mass to enter the pair-instability regime (Spera & Mapelli 2017; Renzo et al. 2020). Namely, stars developing He cores in the 64–135 M range at the end of carbon burning eventually explode as (pair-instability) supernovae leaving no compact remnant (Heger & Woosley 2002; Woosley et al. 2007; Yungelson et al. 2008), whereas stars that develop even more massive He cores efficiently undergo photodisintegration and possibly collapse into IMBHs (Spera & Mapelli 2017). Also, the retention of the envelope can trigger a dredge-up phase during the core helium-burning phase, remove mass from the stellar core and, in turn, prevent pair-production episodes (Costa et al. 2021).
@@
Here, we investigate the formation of IMBHs from VMSs. We account for optically thick winds to assess how different metallicities and mass-loss rates affect the formation of these objects. To this purpose, we have run an extensive set of VMS models with mesa (Paxton et al. 2011, 2018) to explore the impact of stellar winds on the mass of the VMS cores and the resulting BH masses. We find that IMBH formation is strongly suppressed by our optically thick wind model, even at the metallicity of the LMC.
This paper is organized as follows. In Sec. 2 we introduce our models for stars and stellar winds. In Sec. 3, we show the BH mass distributions from single and binary stars in presence of optically thick winds. We summarize our results and conclusion in Sec. 4.
@@ 2 Methods
We model the evolution of VMSs with the hydro-static stellar evolution code mesa (version r12115; Paxton et al. 2011, 2013, 2015, 2018, 2019; Jermyn et al. 2023) adopting the radiation-driven wind models by S23. First, we evolve VMSs until the last phases before core collapse. Then, we use the population-synthesis code sevn (Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023) to explore the BH mass distribution of single and binary VMSs with different wind models and to study a population of massive binary stars. In the following, we introduce our initial conditions and model parameters.
Figure 1: Masses of the He (circles) and CO cores (diamonds) for the wind model from S23 (full markers) and C15 (void markers). The dark (light) gray shaded areas represent the regime for (pulsational) pair-instability supernovae.
@@ 2.1 VMS models with mesa
We adopt the same initial set-up as Simonato et al. (2025), based on the wind mass-loss scheme by S22 and S23111The mesa inlists and run_star_extras files by Sabhahit et al. (2023) are publicly available at this link.. This wind model is based on the concept of transition mass loss introduced by Vink & Gräfener (2012): the transition from optically thin to optically thick winds takes place when the Eddington ratio of the star ($\Gamma_{\mathrm{Edd}}$) exceeds the ratio at the transition point $\Gamma_{\mathrm{Edd,tr}}$. At each iteration, we calculate the mass-loss rate $\dot{M}_{\mathrm{Vink}}$, the escape velocity $v_{\mathrm{esc}}$ and the terminal speed $v_{\infty}$ from the stellar luminosity as described in S23. We use these quantities to calculate the wind efficiency parameter, which determines when the switch occurs:
@@
where $\tau_{F,s}$ is the the flux-weighted mean optical depth at the sonic point. In the optically thick regime, the wind mass-loss follows Vink et al. (2011):
@@
where $\dot{M}_{\mathrm{tr}}$, $L_{\mathrm{tr}}$ and $M_{\mathrm{tr}}$ are the mass-loss rate, the luminosity, and the mass at the transition point, respectively. During the main sequence, we use Eq. 2 to describe the wind mass-loss in the optically thick regime. Otherwise, we refer to Vink et al. (2001). In the core He-burning phase, we use Eq. 2 for temperatures between $4\times 10^{3}\,\mathrm{K}$ and $10^{5}$ K, while for $T<4\times 10^{3}\,\mathrm{K}$, we use the formalism by de Jager et al. (1988). For $T>10^{5}$ K and central hydrogen mass fraction $X_{\text{C}}<0.01$, we adopt the Wolf-Rayet mass-loss by Sander & Vink (2020).
We model convective mixing with the mixing length theory (Cox & Giuli 1968), adopting a constant mixing-length ratio $\alpha_{\text{MLT}}=1.5$. We determine the convective boundaries by applying the Ledoux criterion (Ledoux 1947). We use the semi-convective diffusion parameter $\alpha_{\text{SC}}=1$. We choose the exponential overshooting formalism by Herwig (2000), with efficiency parameter $f_{\text{OV}}=0.03$, to describe this process above the convective regions.
Finally, we activate the co_burn net (parameter auto_extend_net = true) after the main sequence to have a complete and extended network for C- and O-burning and $\alpha$-chains. Following Simonato et al. (2025), we define the core radius of each element as the radius inside which the fraction of the lighter element below $f_{\rm X}=10^{-4}$. For example, the helium (carbon) core radius is the radius inside which the mass fraction of H (4He) drops below $f_{\rm X}=10^{-4}$ ($f_{\rm Y}=10^{-4}$). Our definition is the same as the MIST libraries, and yields an accurate description of the core boundary (Paxton et al. 2011, 2013, 2015; Dotter 2016; Choi et al. 2016).
@@ 2.2 Initial conditions and stopping criteria
We consider a grid of VMS models with masses from 50 to 500 M, by intervals of 25 M. We consider 14 metallicities from $Z=10^{-4}$ to $Z=0.02$. The initial He mass fraction is $Y=Y_{\text{prim}}+Z\,\Delta Y/\Delta Z$, with $Y_{\text{prim}}=0.24$ the primordial He abundance and $\Delta Y/\Delta Z$ is parametrized so that we range from a primordial ($Y=0.24$, $Z=0$) to nearly a solar abundance ($Y=0.28$, $Z=0.02$, Pols et al. 1998). Finally, we calculate the H mass fraction as X=1YZX=1-Y-Z. We consider only non-rotating stellar models.
We evolve the mesa models until the star exceeds a central temperature $\log{T_{\mathrm{c}}/\mathrm{K}}=9.55$, corresponding to the very last stages before core collapse, or it undergoes (pulsational) pair-instability. In the latter case, the creation of electron-positron pairs removes thermal pressure from the gas, possibly triggering a state of global dynamical instability. Because we use the hydro-static integrator of mesa, we cannot integrate the models until (pulsational) pair instability kicks in, but we verify a posteriori the global stability of the star. Specifically, we calculate the first adiabatic exponent $\Gamma_{1}$ (e.g., see also Marchant et al. 2019; Farmer et al. 2019, 2020; Costa et al. 2021), properly weighted and integrated over the mass domain (Stothers 1999):
@@
Figure 2: Final BH mass as a function of $M_{\mathrm{ZAMS}}$ for different metallicities $Z$, for the S23 (upper) and C15 (lower) wind mass-loss prescription. The grey shaded area shows the resulting pair-instability mass gap for the two wind models.
@@ 2.3 BH masses
If the star enters the (pulsational) pair-instability regime, we calculate the final BH mass with the fitting formulas from Mapelli et al. (2020), based on the models by Woosley (2017). In this model, a star undergoes a pulsational pair-instability supernova if the He core has a final mass in the range $32\text{ M}_{\odot}\leq M_{\mathrm{He,f}}<64\text{ M}_{\odot}$. The mass of the BH is therefore estimated as
@@
where $M_{\mathrm{CCSN}}$ is the expected BH mass after a core-collapse supernova, based on the fitting formulas by Fryer et al. (2012), and $\alpha_{\mathrm{P}}$ depends on the mass of He-core and on the ratio between the He-core mass and the total mass in the pre-supernova stage.
@@
Figure 3: Contour plot showing the final BH mass as a function of $M_{\mathrm{ZAMS}}$ and $Z$, for the S23 (upper panel) and the C15 (lower panel) models. The white contours highlight the levels at $20\,{\rm M}_{\odot}$ and $100\,{\rm M}_{\odot}$.
Mass ejection can also occur in stars that undergo a failed supernova because of the shock triggered by instantaneous neutrino loss (Lovegrove & Woosley 2013; Fernández et al. 2018). The ejected mass depends on the ejected energy and the compactness of the star (Fernández et al. 2018). In particular, the ejected energy is a decreasing function of the core compactness, $\xi_{\mathrm{2.5}}$ (O’Connor & Ott 2011). Also, the ejected mass decreases with the stellar envelope compactness, $\xi_{\mathrm{env}}=M_{*}/R_{*}$, where $R_{*}$ is the total radius of the star. The ejected mass is a decreasing fraction of $\xi_{\mathrm{env}}$ (Fernández et al. 2018), and ranges from $M_{\mathrm{ej}}\sim 10\,{\rm M}_{\odot}$ for red supergiant stars ($\xi_{\mathrm{env}}\sim 10^{-2}$) to $M_{\mathrm{ej}}\lesssim 10^{-3}\,{\rm M}_{\odot}$ if the star dies as a Wolf-Rayet object ($\xi_{\mathrm{env}}\sim 10$).
Following Costa et al. (2022a), we estimate the BH mass for the stars that do not undergo a (pulsational) pair instability supernova as:
@@
where $M_{*}$ is the final stellar mass, $\delta M_{\mathrm{G}}$ is the instantaneous gravitational mass loss, and $M_{\mathrm{ej}}$ is the the mass induced by neutrino-driven shocks in a failed supernova. Here, we assume $\delta M_{\mathrm{G}}=0.3\,M_{\odot}$ (Fernández et al. 2018; Costa et al. 2025), which typically yields a corresponding ejected energy between $10^{47}$ and $5\times 10^{47}$ erg. We thus calculate $M_{\mathrm{ej}}$ as the mass of the star with a binding energy less than $3\times 10^{47}$ erg.

### v2

@@
   Giuliano Iorio[Uncaptioned image]    Erika Korb[Uncaptioned image]
@@
Intermediate-mass black holes (IMBHs) are the link between stellar-mass and supermassive black holes. Gravitational waves have started unveiling a population of IMBHs in the $\sim 100-300\,{\rm M}_{\odot}$ range. Here, we investigate the formation of IMBHs from non-rotating very massive stars (VMSs, $100\,{}{\rm M}_{\odot}$). We calculate new VMS models that account for the transition from optically thin to optically thick winds, and study how this enhanced mass loss affects IMBH formation and the black hole mass function at intermediate and high metallicity ($Z=10^{-4}-0.02$). We show that optically thick winds suppress the formation of IMBHs from direct VMS collapse at metallicities $Z>0.001$, one order of magnitude lower than predicted by previous models. Our models indicate that the stellar progenitors of GW231123 must have had a metallicity $Z<0.002$, if the primary black hole formed via direct VMS collapse.
@@ 1 Introduction
Intermediate-mass black holes (IMBHs), with masses $M_{\mathrm{BH}}\sim{10^{2}-10^{5}}$ M are the link between stellar-mass and supermassive black holes at the center of galaxies (greene2020). In the last few years, gravitational waves have provided the first compelling evidence of their existence. The merger remnant of GW190521 represents the first IMBH ever detected via gravitational waves, with a mass of $142^{+28}_{-16}\,\mathrm{M_{\odot}}$ (abbottGW190521; abbottGW190521astro). More recently, parameter estimation for the event GW231123 (gw231123) reveals a black hole (BH) remnant with $m\sim 240\,{\rm M}_{\odot}$. Both BBH components of GW231123 have masses within or above the pair-instability mass gap (Heger2002; Woosley2007), with peculiarly high spins ($\chi\gtrsim 0.7$) that hint at the dynamical origin of this merger (gerosa2017; fishbach2017; antonini2019; gerosa2021; mapelli2021; antonini2025; kiroglu2025b; kiroglu2025). The current fourth gravitational-wave transient catalog (GWTC4pop; GWTC4) already contains a number of other IMBH candidates, which is expected to increase with the next observing runs.
@@
At the same time, searches for IMBHs in the Milky Way globular clusters have mostly led to inconclusive results (mezcua2017), with only a few debated candidate BHs (millerjones2012; tremou2018; baumgardt2019). One notable exception is represented by $\omega$ Centauri, which hosts an IMBH with mass $\geq{}8\times 10^{3}\,{\rm M}_{\odot}$, as revealed by proper motions based on over 20 years of HST data (haberle2024).
@@
Possible IMBH formation pathways include the direct collapse of very massive stars (VMS), with masses above 100 M (Belkus2007; Crowther2010; Bestenlenher2014; vink2015; costa2025; shepherd2025), stellar collisions in dense young star clusters (Portegies2002; freitag2006a; freitag2006b; vanbeveren2009; mapelli2016; dicarlo2021; torniamenti2022), and hierarchical BH mergers in dense and massive star clusters (antonini2019; antonini2023; mapelli2021; rizzuto2021; arcasedda2023; torniamenti2024).
@@
VMSs have been observed in star forming regions and young massive clusters (vink2015), like the Arches cluster near the Galactic Center (martins2008), NGC 3603 (Crowther2010), and R136 in the Large Magellanic Cloud (LMC, Crowther2010; Bestenlenher2014; bestenlehner2020). The estimated initial masses of these objects are up to $\sim 300\,{\rm M}_{\odot}$ (Crowther2010). Also, most of them are observed as single stars (Crowther2010), which may hint that their formation is triggered or enhanced by stellar collisions during the earliest stellar phases (portegieszwart1999; mapelli2016). The possibility that VMSs form IMBHs, however, strongly depends on the amount of mass they lose during their life (vink2018).
@@
At solar metallicity, VMSs can lose a large fraction of their initial mass already during the main sequence, and leave relatively low-mass compact objects (<30<30 M, belczynski2010; mapelli2013; romagnolo2024). At low metallicity, these stars could preserve enough mass to enter the pair-instability regime (Spera2017; renzo2020). Namely, stars developing He cores in the 64–135 M range at the end of carbon burning eventually explode as (pair-instability) supernovae leaving no compact remnant (Heger2002; Woosley2007; yungelson2008), whereas stars that develop even more massive He cores efficiently undergo photodisintegration and possibly collapse into IMBHs (Spera2017). Also, the retention of the envelope can trigger a dredge-up phase during the core helium-burning phase, remove mass from the stellar core and, in turn, prevent pair-production episodes (Costa2021).
@@
Here, we investigate the formation of IMBHs from VMSs. We account for optically thick winds to assess how different metallicities and mass-loss rates affect the formation of these objects. To this purpose, we have run an extensive set of VMS models with mesa (Paxton2011; Paxton2018) to explore the impact of stellar winds on the mass of the VMS cores and the resulting BH masses. We find that IMBH formation is strongly suppressed by our optically thick wind model, even at the metallicity of the LMC.
This paper is organized as follows. In Sec. 2 we introduce our models for stars and stellar winds. In Sec. 3, we show the BH mass distributions from single and binary stars in presence of optically thick winds. In Sect. 4 we discuss the possible impact of rotation and core overshooting. Finally we summarize our results and conclusion in Sec. 5.
@@ 2 Methods
We model the evolution of VMSs with the hydro-static stellar evolution code mesa (version r12115; Paxton2011; Paxton2013; Paxton2015; Paxton2018; Paxton2019; jermy2023) adopting the radiation-driven wind models by Sabhahit2023. First, we evolve VMSs until the last phases before core collapse. Then, we use the population-synthesis code sevn (Spera2019; Mapelli2020; Iorio2023) to explore the BH mass distribution of single and binary VMSs with different wind models and to study a population of massive binary stars. In the following, we introduce our initial conditions and model parameters.
Figure 1: Masses of the He (circles) and CO cores (diamonds) for the wind model from Sabhahit2023 (full markers) and Chen2015 (void markers). The dark (light) gray shaded areas represent the regime for (pulsational) pair-instability supernovae.
@@ 2.1 VMS models with mesa
We adopt the same initial set-up as simonato2025, based on the wind mass-loss scheme by Sabhahit2022 and Sabhahit2023111The mesa inlists and run_star_extras files by Sabhahit2023 are publicly available at this link.. This wind model is based on the concept of transition mass loss introduced by Vink2012: the transition from optically thin to optically thick winds takes place when the Eddington ratio of the star ($\Gamma_{\mathrm{Edd}}$) exceeds the ratio at the transition point $\Gamma_{\mathrm{Edd,tr}}$. At each iteration, we calculate the mass-loss rate $\dot{M}_{\mathrm{Vink}}$, the escape velocity $v_{\mathrm{esc}}$ and the terminal speed $v_{\infty}$ from the stellar luminosity as described in Sabhahit2023. We use these quantities to calculate the wind efficiency parameter, which determines when the switch occurs:
@@
where $\tau_{F,s}$ is the the flux-weighted mean optical depth at the sonic point. In the optically thick regime, the wind mass-loss follows Vink2011:
@@
where $\dot{M}_{\mathrm{tr}}$, $L_{\mathrm{tr}}$ and $M_{\mathrm{tr}}$ are the mass-loss rate, the luminosity, and the mass at the transition point, respectively. During the main sequence, we use Eq. 2 to describe the wind mass-loss in the optically thick regime. Otherwise, we refer to Vink2001. In the core He-burning phase, we use Eq. 2 for temperatures between $4\times 10^{3}\,\mathrm{K}$ and $10^{5}$ K, while for $T<4\times 10^{3}\,\mathrm{K}$, we use the formalism by deJager1988. For $T>10^{5}$ K and central hydrogen mass fraction $X_{\text{C}}<0.01$, we adopt the Wolf-Rayet mass-loss by Sander2020.
We model convective mixing with the mixing length theory (Cox1968), adopting a constant mixing-length ratio $\alpha_{\text{MLT}}=1.5$. We determine the convective boundaries by applying the Ledoux criterion (Ledoux1947). We use the semi-convective diffusion parameter $\alpha_{\text{SC}}=1$. We choose the exponential overshooting formalism by Herwig2000, with efficiency parameter $f_{\text{OV}}=0.03$, to describe this process above the convective regions.
Finally, we activate the co_burn net (parameter auto_extend_net = true) after the main sequence to have a complete and extended network for C- and O-burning and $\alpha$-chains. Following simonato2025, we define the core radius of each element as the radius inside which the fraction of the lighter element below $f_{\rm X}=10^{-4}$. For example, the helium (carbon) core radius is the radius inside which the mass fraction of H (4He) drops below $f_{\rm X}=10^{-4}$ ($f_{\rm Y}=10^{-4}$). Our definition is the same as the MIST libraries, and yields an accurate description of the core boundary (Paxton2011; Paxton2013; Paxton2015; Dotter2016; Choi2016).
@@ 2.2 Initial conditions and stopping criteria
We consider a grid of VMS models with masses from 50 to 500 M, by intervals of 25 M. We consider 14 metallicities from $Z=10^{-4}$ to $Z=0.02$. The initial He mass fraction is $Y=Y_{\text{prim}}+Z\,\Delta Y/\Delta Z$, with $Y_{\text{prim}}=0.24$ the primordial He abundance and $\Delta Y/\Delta Z$ is parametrized so that we range from a primordial ($Y=0.24$, $Z=0$) to nearly a solar abundance ($Y=0.28$, $Z=0.02$, Pols1998) [^2] Here, we refer to the traditionally considered value for solar metallicity $Z=0.02$ (grevesse1998; but see, e.g., magg2022 for more recent estimates).. Finally, we calculate the H mass fraction as X=1YZX=1-Y-Z. We consider only non-rotating stellar models.
We evolve the mesa models until the star exceeds a central temperature $\log{T_{\mathrm{c}}/\mathrm{K}}=9.55$, corresponding to the very last stages before core collapse, or it undergoes (pulsational) pair-instability. In the latter case, the creation of electron-positron pairs removes thermal pressure from the gas, possibly triggering a state of global dynamical instability. Because we use the hydro-static integrator of mesa, we cannot integrate the models until (pulsational) pair instability kicks in, but we verify a posteriori the global stability of the star. Specifically, we calculate the first adiabatic exponent $\Gamma_{1}$ (e.g., see also Marchant2019; farmer19; Farmer2020; Costa2021), properly weighted and integrated over the mass domain (stothers1999):
@@
Figure 2: Final BH mass as a function of $M_{\mathrm{ZAMS}}$ for different metallicities $Z$, for the Sabhahit2023 (upper) and Chen2015 (lower) wind mass-loss prescription. The grey shaded area shows the resulting pair-instability mass gap for the two wind models.
@@ 2.3 BH masses
If the star enters the (pulsational) pair-instability regime, we calculate the final BH mass with the fitting formulas from Mapelli2020, based on the models by Woosley2017. In this model, a star undergoes a pulsational pair-instability supernova if the He core has a final mass in the range $32\text{ M}_{\odot}\leq M_{\mathrm{He,f}}<64\text{ M}_{\odot}$. The mass of the BH is therefore estimated as
@@
where $M_{\mathrm{CCSN}}$ is the expected BH mass after a core-collapse supernova, based on the fitting formulas by Fryer2012, and $\alpha_{\mathrm{P}}$ depends on the mass of He-core and on the ratio between the He-core mass and the total mass in the pre-supernova stage.
@@
Figure 3: Contour plot showing the final BH mass as a function of $M_{\mathrm{ZAMS}}$ and $Z$, for the Sabhahit2023 (upper panel) and the Chen2015 (lower panel) models. The white contours highlight the levels at $20\,{\rm M}_{\odot}$ and $100\,{\rm M}_{\odot}$.
Mass ejection can also occur in stars that undergo a failed supernova because of the shock triggered by instantaneous neutrino loss (lovegrove2013; fernandez2018). The ejected mass depends on the ejected energy and the compactness of the star (fernandez2018). In particular, the ejected energy depends on the core compactness, $\xi_{\mathrm{2.5}}$ (oconnor2011; burrows2021; burrows2024). Also, the ejected mass decreases with the stellar envelope compactness, $\xi_{\mathrm{env}}=M_{*}/R_{*}$, where $R_{*}$ is the total radius of the star. The ejected mass depends on $\xi_{\mathrm{env}}$ (fernandez2018), and ranges from $M_{\mathrm{ej}}\sim 10\,{\rm M}_{\odot}$ for red supergiant stars ($\xi_{\mathrm{env}}\sim 10^{-2}$) to $M_{\mathrm{ej}}\lesssim 10^{-3}\,{\rm M}_{\odot}$ if the star dies as a Wolf-Rayet object ($\xi_{\mathrm{env}}\sim 10$).
Following Costa2022, we estimate the BH mass for the stars that do not undergo a (pulsational) pair instability supernova as:
@@
where $M_{*}$ is the final stellar mass, $\delta M_{\mathrm{G}}$ is the instantaneous gravitational mass loss, and $M_{\mathrm{ej}}$ is the the mass induced by neutrino-driven shocks in a failed supernova. Here, we assume $\delta M_{\mathrm{G}}=0.3\,M_{\odot}$ (fernandez2018; costa2025), which typically yields a corresponding ejected energy between $10^{47}$ and $5\times 10^{47}$ erg. We thus calculate $M_{\mathrm{ej}}$ as the mass of the star with a binding energy less than $3\times 10^{47}$ erg.